# Estimating the Influence of Sequentially Correlated Literary Properties in Textual Classification: A Data-Centric Hypothesis-Testing Approach

### v2

@@ 3 Hypothesis-Testing Scheme
$MCC=\frac{TN\times TP-FN\times FP}{\sqrt{(TP+FP)(TP+FN)(FN+FP)(TN+FN)}},$ (4)
where T (F) stands for True (False) and P (N) stands for Positive (Negative), which are evaluations of some sequence against another. We use MCC to evaluate the agreement between the reference label sequence $L^{(\rm{real})}$ and $L$, or a drawn label sequence $L^{(null)}$ as our statistic of choice, normalized to percent within the range of 50%–100%, where 50% suggests an arbitrary overlap between two label sequences, and 100% suggests perfect overlap. By drawing multiple random correlated label sequences $L^{(\rm{null})}$, we empirically estimate the null distribution of MCC between the random and $L^{(\rm{real})}$. We can derive the probability of the MCC score between the latter and $L$ to be drawn from that distribution – implying that the unsupervised partition is based predominantly on sequentially correlated literary properties.
We present the procedure of generating a null label sequence $L^{\rm{(null)}}$ using this method schematically in Algorithm 1.
Algorithm 1 Hypothesis-testing Scheme: Generating a null label sequence $L^{\rm{(null)}}$
@@
6:Impose elementwise thresholding on $X$, such that $X_{i}=1\Leftrightarrow X_{i}>0.5$, and $X_{i}=0$ otherwise.
7:$X$ is now $L^{\rm{\rm{(null)}}}$.
@@ 4.1 Rationale
Our research is targeted primarily to assist in the analysis of historical texts, which are often multilayered and redacted. Unlike the analysis of modern texts (whose authors are known) predominantly used to explore the underpinnings of linguistic patterns and structure in literary works, studying ancient texts often serves a different purpose. Specifically, tracing the text’s origin to its extant version often requires interdisciplinary efforts across linguistics, history, and computer science disciplines (e.g., Bühler et al., 2024).
Given the nature of such endeavors, where insights from multiple fields converge, it is imperative to employ interpretable methods. Historical texts pose unique challenges that demand a heightened focus on interpretability in analysis (Chapman, 2017). They often contain layers of annotations, marginalia, and modifications made by scribes or editors over time, obscuring the original authorial intent (e.g., Müller et al., 2014). Additionally, historical texts may be written in archaic or obsolete languages, further complicating automated analysis (e.g., Baum, 2017). Moreover, the historical context surrounding these texts, including cultural norms, linguistic conventions, and socio-political factors, must be carefully considered to ensure accurate interpretation.
Given these complexities, employing interpretable methods is essential for unraveling the intricate layers of meaning embedded within historical texts (Chapman, 2017). This approach enables researchers to navigate the nuances of language and context for a deeper understanding of historical narratives and societal dynamics.
Therefore, rather than solely relying on state-of-the-art neural language models, we emphasize the exploration of ”older” features that have proven efficacy in textual analysis. By prioritizing interpretability and drawing on established methodologies, our approach ensures that the insights obtained from our study are accessible and meaningful across disciplinary boundaries. This commitment to methodological transparency and interpretability strengthens the interdisciplinary dialogue and enhances the applicability of our research findings.
That said, by leveraging traditional textual analysis techniques alongside modern neural embeddings, we aim to discern how well each method captures the unique stylistic elements. Through supervised experiments, we train models to differentiate between pairs of texts that were either composed by the same author or not. Specifically, we wish to explore potential biases inherent in supervised approaches, as they may rely on labeled data that could inadvertently reflect the perspectives or assumptions of annotators.
In contrast, unsupervised experiments allow for the exploration of the underlying structure of texts without predefined annotation. By applying clustering algorithms to traditional and neural embeddings, we can uncover patterns of stylistic variation and similarity within our dataset. This approach enables us to examine whether neural embeddings offer a more nuanced representation of stylistic features, thereby facilitating a deeper understanding of the authorship of texts.
Through these experiments, coupled with our hypothesis-testing approach, we seek to demonstrate the strengths and limitations of both traditional and neural methods in capturing the stylistic components of texts and the susceptibility of supervised approaches to training bias under different embedding setups.
Lastly, we aim to demonstrate that our hypothesis-testing approach provides critical insights into the basis of textual classification results, substantially enhancing interpretability. By quantifying the extent to which classification depends on sequentially versus non-sequentially correlated features, our method enables a nuanced understanding of the factors driving distinctions between texts. This deeper analysis clarifies when specific elements, such as thematic content, rather than authorial style, play a primary role, allowing researchers to discern the true drivers of classification outcomes.

### v4

@@ 3 Hypothesis-Testing Scheme
$\text{MCC}=\frac{\text{TN}\times\text{TP}-\text{FN}\times\text{FP}}{\sqrt{(\text{TP}+\text{FP})(\text{TP}+\text{FN})(\text{FN}+\text{FP})(\text{TN}+\text{FN})}},$ (4)
where T (F) stands for true (false) and P (N) stands for positive (negative), which are evaluations of some sequence against another. We use MCC to evaluate the agreement between the reference label sequence $L^{(\mathrm{real})}$ and $L$, or a drawn label sequence $L^{(null)}$ as our statistic of choice, normalized to percent within the range of 50%–100%, where 50% suggests an arbitrary overlap between two label sequences, and 100% suggests perfect overlap. By drawing multiple random correlated label sequences $L^{(\mathrm{null})}$, we empirically estimate the null distribution of MCC between the random and $L^{(\mathrm{real})}$. We can derive the probability of the MCC score between the latter and $L$ to be drawn from that distribution – implying that the unsupervised partition is based predominantly on sequentially correlated literary properties.
We present the procedure of generating a null label sequence $L^{\mathrm{(null)}}$ using this method schematically in Algorithm 1.
Algorithm 1 Hypothesis-Testing Scheme: Generating a null label sequence $L^{\mathrm{(null)}}$
@@
6:Impose element-wise thresholding on $X$, such that $X_{i}=1\Leftrightarrow X_{i}>0.5$, and $X_{i}=0$ otherwise.
7:$X$ is now $L^{\mathrm{\mathrm{(null)}}}$.
@@ 4.1 Rationale
Rather than relying solely on state-of-the-art neural language models, we emphasize the value of traditional features that have demonstrated long-standing efficacy in textual analysis. This is particularly relevant for unannotated historical texts, which are often redacted or multilayered, and present stylistic and structural complexities not easily captured by modern embeddings (e.g., Bühler et al., 2024). Frequency-based representations such as tf-idf over word or character $n$-grams remain powerful tools in such contexts.
To complement these, we incorporate a Delta-like embedding scheme inspired by Burrows’ Delta (Burrows, 2002), which uses Z-score standardized frequencies of the most frequent $n$-grams or $k$-mers to construct interpretable, content-agnostic stylistic profiles. This representation enhances interpretability and serves as a valuable baseline for clustering and comparison, aligning our pipeline with stylometric tradition.
Our aim is to evaluate how effectively various embeddings capture non-sequentially correlated literary properties. In supervised experiments, we train models to differentiate between text pairs authored either by the same or different individuals. This allows us to examine how supervised systems may encode implicit biases from training labels, potentially reflecting thematic or structural assumptions.
Unsupervised experiments, by contrast, reveal the intrinsic structure of stylistic variation without relying on annotation. By applying clustering to traditional, Delta-style, and neural embeddings, we probe whether modern embeddings offer finer stylistic discrimination or whether classic stylometric indicators remain competitive, especially in genre-diverse or data-scarce contexts.
To further enhance interpretability, we apply a hypothesis-testing framework that quantifies the degree to which classifications rely on sequentially versus non-sequentially correlated features. This enables us to distinguish between classifications driven by authorial style and those influenced by thematic content, offering a deeper understanding of model behavior and interpretive validity.
@@ 4.2 Corpora
We conduct our analysis on several curated English prose corpora selected to represent a broad range of literary styles, genres, and narrative structures. This diversity is designed to stress-test our hypothesis-testing framework under realistic stylometric challenges and to reflect both classical and contemporary authorial voices.